# Characterization of non-Markovianity with maximal extractable qubit-reservoir entanglement

### v1

@@ Abstract
When the spectral structure of the reservoir is modified, it can become non-Markovian. Decaying resonators represent the most typical type of non-Markovian reservoirs, which have a Lorentzian spectrum 1 ; 2 . The memory effect of the reservoir depends on the ratio of the system-reservoir coupling to the spectral linewidth. When this ratio tends to infinity, the spectral spread can be neglected, so that the reservoir reduces to a single-mode resonator, which exchanges energy with the system in a reversible manner. In the opposite limit, the spectral spread can be considered infinite, so that the reservoir is approximately Markovian. In the intermediate regime, the reservoir displays some memory effect, which is gradually lost with the decrease of the coupling strength.
@@
In the past decade, many efforts have been devoted to the quantification of non-Markovianity 6 ; 7 ; 8 ; 9 ; 10 . It has been shown that the quantum memory effect can be quantified by tracking the evolutions of two initially distinct quantum states in the reservoir 6 . When the reservoir is Markovian, the distinguishability between these two states is monotonously reduced and finally tends to zero when the information is completely leaked into the environment. For a non-Markovian process, this distinguishability displays an oscillatory pattern, as a consequence of the reversed information flow from the reservoir back to the system. Alternatively, the non-Markovianity can also be measured by using an ancilla that is initially entangled with the system in the reservoir, and tracking the ancilla-system entanglement during the system-reservoir interaction 7 . The reservoir’s memory effect is characterized by the extent to which the lost entanglement can be revived. Based on these measures, the non-Markovian effects have been experimentally explored in different systems, including photonic qubits 11 ; 12 ; 13 , nitrogen-vacancy centers 14 ; 15 ; 16 , and ion traps 17 . Recently, Gaikwad et al. explored the crossover between the Markovian and non-Markovian dynamics with a superconducting processor 18 , where the entanglement between a test qubit and an ancilla is used as a measure to quantify the non-Markovianity of the environment of the test qubit, comprised of a third qubit and its readout resonator.
@@
We here propose an alternative measure for quantifying the memory effect for a general non-Markovian reservoir that interacts with a qubit, and present an experimental demonstration. Instead of introducing an ancilla that is initially entangled with the test qubit, the non-Markovianity is directly quantified by the extractable entanglement between the qubit and the reservoir generated by their interaction. When the qubit is initially in the excited state, the excitation is shared by the system and the reservoir due to their interaction. This excitation sharing results in the qubit-reservoir entanglement. The amount of the extractable maximum entanglement depends the non-Markovianity of the reservoir. In the strong coupling regime, the system-reservoir dynamics approximately reduces to the vacuum Rabi oscillation of the Jaynes-Cummings (JC) model, where the qubit-photon entanglement is completely extractable, with the maximum value approximating 1. With the decrease of the coupling strength, the amount of the extractable maximum entanglement is gradually reduced due to the spreading of the spectrum. In the Markovian limit, no entanglement can be extracted. The experimental data used to demonstrate the crossover between the Markovian and non-Markovian dynamics is taken from a previous experiment 19 , where a non-Hermitian JC model was synthesized with a superconducting qubit controllably coupled to its readout resonator. Here the readout resonator is taken as a reservoir. The experimental results confirm the theoretical predictions.
@@
$H=\left|g\right\rangle\left\langle e\right|\displaystyle\int\limits{}_{0}^{\infty}d\omega J(\omega)\lambda(\omega)e^{i\delta(\omega)t}a^{\dagger}(\omega)+H.c.,$ (1)
@@
When the photon is leaked into the Markovian reservoir, it cannot be re-absorbed by the qubit and the information is completely lost. Once this event occurs, both $Q$ and $M_{0}$ jump to their ground state. The correlation between the Markovian reservoir and the qubit or the bosonic mode cannot be extracted. The part of quantum correlation that can be extracted is that associated with the non-unitarily evolved wavefunction $\left|\psi(t)\right\rangle$. This entanglement is characterized by the concurrence 20
@@
The Concurrence ${\cal C}$ measures the unconditional entanglement between $Q$ and $M_{0}$, averaged over the trajectories with and without quantum jumps.
Figure 1: (color online). Theoretical results. a Extractable concurrence (${\cal C}$) between a qubit and a reservoir of photonic modes versus $\xi$ and $\tau$. The reservoir, initially in the vaccum state, is composed of photonic modes with a Lorentzian spectrum with a line width $\kappa$. The qubit, initially in the excited state, is on resonance with the central frequency of the continuum, $\omega_{0}$. Here $\tau=\kappa t/4$ denotes the rescaled interaction time, and $\xi$ is the ratio of the effective coupling strength between the qubit and the photonic mode with frequency $\omega_{0}$ to $\kappa/4$, and $\tau$. b Maximum of $\mathcal{C}$ during the evolution versus $\xi$. The dashed line is the derivative of $\mathcal{C}_{\max}$ with respect to $\xi$.
@@
To confirm the validity of the approach, we use the data measured in a previous experiment focusing on the exceptional entanglement phase transition in a non-Hermitian qubit-photon model 19 . The experiment was performed with a circuit electrodynamics system, where an Xmon qubit with a negligible dissipative rate was coupled to its readout resonator with a non-negligible photonic decaying rate $\kappa=5$ MHz through a sideband interaction, mediated by a parametric modulation. Here the readout resonator is taken as the reservoir for the qubit, other than a part of the system. The effective qubit-photon coupling strength is controllable via the modulating amplitude. In the previous work we display the evolution of the qubit-photon entanglement associated with the no-jump trajectory, calculated with the output qubit-photon state projected onto the single-excitation subspace $\left\{\left|e,0\right\rangle,\left|g,1\right\rangle\right\}$. Such a conditional output state was obtained by reconstructing the qubit-photon density matrix with the assistance of an ancilla qubit, and then discarding the element associated to $\left|g,0\right\rangle$. The evolution of the post-selected output state coincides with the no-jump trajectory governed by the non-Hermitian Hamiltonian. On the contrary, as discussed above, the non-Markovianity is manifested here by the entanglement evolution associated with the whole output density matrix, reconstructed in the subspace $\left\{\left|e,0\right\rangle,\left|g,1\right\rangle,\left|g,0\right\rangle\right\}$.
Figure 2: (color online). Experimental results. a Measured ${\cal C}$ versus $\xi$ and $\tau$. The data is taken from a previous experiment 19 , which was performed in a circuit QED system, where the readout resonator of an Xmon qubit with a photonic decaying rate $\kappa$ is taken as the reservoir. b Measured ${\cal C}_{\max}$ versus $\xi$.
@@
In conclusion, we have shown that the maximal entanglement between a qubit and its reservoir extractable an be used as a measure for quantifying the non-Markovianity of the dynamics. When the qubit-reservoir coupling is much smaller than the spectral line-width of the reservoir, the leakage of the excitation from the qubit to the reservoir is almost irreversible. Consequently, most of the qubit-reservoir entanglement produced by their interaction is unextractable, reflecting the Markovian feature. With the increase of the coupling, the reservoir displays a non-Markovian effect, manifested by the nonvanishing qubit-reservoir entanglement that can be extracted. We illustrate this measure with a circuit QED experiment, where the readout resonator of a qubit acts as a reservoir of photonic modes with a continuous spectrum. By tuning the effective qubit-resonator coupling, we observed the progressive crossover from the Markovian to non-Markovian dynamics.
This work was supported by the National Natural Science Foundation of China (Grant No. 12274080, No. 11875108, No. 12174058, No. 12204105, No. 12374479), the NSF (Grant No. 2329027) and Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0300200).

### v2

@@ Abstract
When the spectral structure of the reservoir is modified, it can become non-Markovian. Due to this spectral modification, the reservoir exhibits a memory effect, which is characterized by the future state of the reservoir depending on its history. This memory effect implies that the information flow from the system to the reservoir becomes partly irreversible. The degree of non-Markovianity is quantified by the extent to which information can flow back from the reservoir to the system. This reversibility is influenced by the ratio between the system-reservoir coupling strength and the spectral width of the reservoir. A higher ratio indicates a greater ability of the reservoir to return information to the system, thereby enhancing the non-Markovian behavior. Decaying resonators represent the most typical type of non-Markovian reservoirs, which have a Lorentzian spectrum 1 ; 2 . The memory effect of the reservoir depends on the ratio of the system-reservoir coupling to the spectral linewidth. When this ratio tends to infinity, the spectral spread can be neglected, so that the reservoir reduces to a single-mode resonator, which exchanges energy with the system in a reversible manner. In the opposite limit, the spectral spread can be considered infinite, so that the reservoir is approximately Markovian. In the intermediate regime, the reservoir displays some memory effect, which is gradually lost with the decrease of the coupling strength.
@@
In the past decade, many efforts have been devoted to the quantification of non-Markovianity 6 ; 7 ; 8 ; 9 ; 10 ; 11 ; 12 ; 13 ; 14 . It has been shown that the quantum memory effect can be quantified by tracking the evolutions of two initially distinct quantum states in the reservoir 6 . When the reservoir is Markovian, the distinguishability between these two states is monotonously reduced and finally tends to zero when the information is completely leaked into the environment. For a non-Markovian process, this distinguishability displays an oscillatory pattern, as a consequence of the reversed information flow from the reservoir back to the system. Alternatively, the non-Markovianity can also be measured by using an ancilla that is initially entangled with the system in the reservoir, and tracking the ancilla-system entanglement during the system-reservoir interaction 7 ; 11 ; 12 ; 13 ; 14 . The non-Markovian degree is quantified by the extent of temporal increase in the entanglement between the open system and the ancilla. The reservoir’s memory effect is characterized by the extent to which the lost entanglement can be revived. Based on these measures, the non-Markovian effects have been experimentally explored in different systems, including photonic qubits 15 ; 16 ; 17 , nitrogen-vacancy centers 18 ; 19 ; 20 , and ion traps 21 . Recently, Gaikwad et al. explored the crossover between the Markovian and non-Markovian dynamics with a superconducting processor 22 , where the entanglement between a test qubit and an ancilla is used as a measure to quantify the non-Markovianity of the environment of the test qubit, comprised of a third qubit and its readout resonator.
@@
We here propose an alternative measure for quantifying the memory effect for a general non-Markovian reservoir that interacts with a qubit, and present an experimental demonstration. Instead of introducing an ancilla that is initially entangled with the test qubit, the non-Markovianity is directly quantified by the extractable entanglement between the qubit and the reservoir generated by their interaction. When the qubit is initially in the excited state, the excitation is shared by the system and the reservoir due to their interaction. This excitation sharing results in the qubit-reservoir entanglement. The amount of the extractable maximum entanglement depends the non-Markovianity of the reservoir. In the strong coupling regime, the system-reservoir dynamics approximately reduces to the vacuum Rabi oscillation of the Jaynes-Cummings (JC) model, where the qubit-photon entanglement is completely extractable, with the maximum value approximating 1. With the decrease of the coupling strength, the amount of the extractable maximum entanglement is gradually reduced due to the spreading of the spectrum. In the Markovian limit, no entanglement can be extracted. The experimental data used to demonstrate the crossover between the Markovian and non-Markovian dynamics is taken from a previous experiment 23 , where a non-Hermitian JC model was synthesized with a superconducting qubit controllably coupled to its readout resonator. Here the readout resonator is taken as a reservoir. The experimental results confirm the theoretical predictions. Compared to previous measures, the present measure offers significant advantages. Notably, it does not require precise comparison between the evolution trajectories of two initially distinct quantum states, as referenced in 6 . Additionally, our method eliminates the need for any ancilla qubit, which has been a requisite in other approaches 7 ; 22 .
@@
[…]
@@
When the photon is leaked into the Markovian reservoir, it cannot be re-absorbed by the qubit and the information is completely lost. Once this event occurs, both $Q$ and $M_{0}$ jump to their ground state. The correlation between the Markovian reservoir and the qubit or the bosonic mode cannot be extracted. The part of quantum correlation that can be extracted is that associated with the non-unitarily evolved wavefunction $\left|\psi(t)\right\rangle$. This entanglement is characterized by the concurrence 24
@@
The concurrence ${\cal C}$ measures the unconditional entanglement between $Q$ and $M_{0}$, averaged over the trajectories with and without quantum jumps. We note that this concurrence quantifies the amount of the qubit-reservoir entanglement that can be extracted. For a Markovian reservoir with $\kappa\rightarrow\infty$, the entanglement between the system and reservoir cannot be extracted, resulting in an extractable system-reservoir entanglement of zero. This implies that such a concurrence acts as a witness of non-Markovianity 11 .
Figure 1: Theoretical results. a Extractable concurrence (${\cal C}$) between a qubit and a reservoir of photonic modes versus $\xi$ and $\tau$. The reservoir, initially in the vaccum state, is composed of photonic modes with a Lorentzian spectrum with a line width $\kappa$. The qubit, initially in the excited state, is on resonance with the central frequency of the continuum, $\omega_{0}$. Here $\tau=\kappa t/4$ denotes the rescaled interaction time, and $\xi$ is the ratio of the effective coupling strength between the qubit and the photonic mode with frequency $\omega_{0}$ to $\kappa/4$, and $\tau$. b Maximum of $\mathcal{C}$ during the evolution versus $\xi$. The dashed line is the derivative of $\mathcal{C}_{\max}$ with respect to $\xi$.
@@
Figure 2: Schematic of the experimental system. The system consists of a Josephson-based artificial atom, known as the Xmon qubit ($Q$), and its associated readout resonator ($R$), which has a decay rate of $\kappa=5$ MHz. This resonator functions as a non-Markovian reservoir for $Q$. The coupling between $Q$ and $R$ is facilitated by a sideband modulation applied parametrically to $Q$.
To confirm the validity of the approach, we use the data measured in a previous experiment focusing on the exceptional entanglement phase transition in a non-Hermitian qubit-photon model 23 . The experiment was performed with a circuit electrodynamics system (Fig. 2), where an Xmon qubit with a negligible dissipative rate was coupled to its readout resonator with a non-negligible photonic decaying rate $\kappa=5$ MHz through a sideband interaction, mediated by a parametric modulation with the amplitude $\varepsilon$ and frequency $\nu$. Here the readout resonator is taken as the reservoir for the qubit, other than a part of the system. When $\nu=(\omega_{r}-\omega_{q})/n$, the qubit is effectively coupled to the resonator at the $n$th upper sideband, with an effective coupling strength $\lambda=gJ_{n}(\varepsilon/\nu)$. Here, $\omega_{q}$ and $\omega_{r}$ denote the frequencies of the qubit and the resonator, respectively, $g$ stands for their on-resonance coupling strength, and $J_{n}(\mu)$ is the $n$th-order Bessel function of the first kind 25 ; 26 ; 27 . In our experiment, the effective qubit-photon coupling strength is carefully controlled by tuning the modulating frequency and amplitude. To minimize the crosstalk, we have employed first- and second-order sideband couplings in the designated regions $10<\xi<$ 1 and $\xi>1$, respectively. This strategic approach ensures that the system operates with the desired precision and reduces unintended interactions. In the previous work we display the evolution of the qubit-photon entanglement associated with the no-jump trajectory, calculated with the output qubit-photon state projected onto the single-excitation subspace $\left\{\left|e,0\right\rangle,\left|g,1\right\rangle\right\}$. Such a conditional output state was obtained by reconstructing the qubit-photon density matrix with the assistance of an ancilla qubit, and then discarding the element associated to $\left|g,0\right\rangle$. The evolution of the post-selected output state coincides with the no-jump trajectory governed by the non-Hermitian Hamiltonian. On the contrary, as discussed above, the non-Markovianity is manifested here by the entanglement evolution associated with the whole output density matrix, reconstructed in the subspace $\left\{\left|e,0\right\rangle,\left|g,1\right\rangle,\left|g,0\right\rangle\right\}$.
@@
Figure 3: Experimental results. a Measured ${\cal C}$ versus $\xi$ and $\tau$. The data is taken from a previous experiment 23 , which was performed in a circuit QED system, where the readout resonator of an Xmon qubit with a photonic decaying rate $\kappa$ is taken as the reservoir. b Measured ${\cal C}_{\max}$ versus $\xi$.
@@
Due to unavoidable experimental imperfections, the measured concurrences are slightly lower than the ideal values depicted in Fig. 1. The primary source of error arises from the influence of neglected off-resonant terms in the expansion of the parametric modulation. These off-resonant terms cause the dynamics of the experimental system to diverge from the idealized model. Specifically, the dominant off-resonant term is the carrier coupling with a strength of $gJ_{0}(\varepsilon/\nu)$, which introduces fast, slight oscillations in the system’s behavior. This effect is consistent with the observation detailed in 23 . However, we note that parametric modulation becomes unnecessary when the resonator frequency falls within the adjustable range of the qubit frequency. In such cases, the unwanted off-resonant terms mentioned earlier are effectively avoided. Additionally, the impact of qubit frequency fluctuations is mitigated by the qubit-resonator interaction, which inherently enhances the system’s robustness against dephasing noise, as explained below. Specifically, the qubit-resonator interaction generates two dressed states separated by an energy gap, thereby preventing dephasing noise from inducing transitions between these dressed states 28 .
In conclusion, we have shown that the maximal entanglement between a qubit and its reservoir extractable can be used as a measure for quantifying the non-Markovianity of the dynamics. When the qubit-reservoir coupling is much smaller than the spectral line-width of the reservoir, the leakage of the excitation from the qubit to the reservoir is almost irreversible. Consequently, most of the qubit-reservoir entanglement produced by their interaction is unextractable, reflecting the Markovian feature. With the increase of the coupling, the reservoir displays a non-Markovian effect, manifested by the nonvanishing qubit-reservoir entanglement that can be extracted. We illustrate this measure with a circuit QED experiment, where the readout resonator of a qubit acts as a reservoir of photonic modes with a continuous spectrum. By tuning the effective qubit-resonator coupling, we observed the progressive crossover from the Markovian to non-Markovian dynamics.
This work was supported by the National Natural Science Foundation of China (Grant Nos. 12274080, 12474356, 12475015, 11875108, 12174058, 12204105, and 12374479), the NSF (Grant No. 2329027) and Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0300200).